# Revisiting the Radial Metallicity Gradient-Age Relation in the Milky Way’s Thin and Thick Disks

### v1

@@ Abstract
Galactic disks typically exhibit a negative radial metallicity gradient, indicating faster enrichment in the inner regions. Recent studies report that this gradient becomes flatter with increasing stellar age in the Milky Way’s (MW) thin disk, while the thick disk exhibits a mildly positive gradient across all ages. In this work, we revisit the metallicity gradient–age relation (MGAR) in both the thin and thick disks of the MW. We use spectroscopic data from LAMOST DR8 and stellar ages calibrated with asteroseismology. Our results show a steadily flattening MGAR in the thin disk and confirm a positive gradient $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$ in the thick disk. The flattening in the thin disk may be caused by large-scale radial migration induced by transient spiral arms, or by a time-dependent steepening of the interstellar medium (ISM) metallicity gradient as suggested by recent FIRE2 simulations. The positive gradient in the thick disk may reflect early enrichment of the outer regions by strong feedback or starburst-driven outflows in a turbulent, gas-rich proto-disk. These findings suggest distinct chemodynamical evolution paths for the MW’s thin and thick disks and provide valuable constraints for future models of Galactic chemical evolution.
Milky Way dynamics (1051) — Milky Way disk (1050) — Milky Way evolution (1052)
@@ 1 Introduction
A cornerstone of galaxy evolution is the negative radial metallicity gradient, which describes the decrease in stellar metallicity with increasing galactocentric radius. Matteucci & Francois (1989) reproduced this trend using “inside-out” formation models, in which the inner disk forms stars more rapidly and enriches earlier than the outer regions. This scenario naturally results in a negative metallicity gradient for stars formed from the interstellar medium (ISM, Matteucci, 2021). Following this framework, Minchev et al. (2018) found that the ISM gradient flattens over time as the outer disk becomes progressively enriched and approaches the central metallicity.
@@
The pioneering study on the MGAR in our MW by Mayor (1976) measured a mean radial metallicity gradient of $0.05\pm 0.01\,\mathrm{dex\,kpc^{-1}}$, which steepens to $0.10\pm 0.02\,\mathrm{dex\,kpc^{-1}}$ for younger stars. Using effective temperature ($T_{\mathrm{eff}}$) of main-sequence stars as a proxy for age, Yu et al. (2012) found that the gradient steepens with increasing $T_{\mathrm{eff}}$ (typically associated with decreasing age), but shows no trend for stars with $T_{\mathrm{eff}}<6,000\,\mathrm{K}$—those with main-sequence lifetimes $\gtrsim 10$ Gyr. These early findings hinted at a flattening gradient for increasingly older stars in the MW’s thin disk.
@@
Large-scale spectroscopic surveys have greatly facilitated precise age estimates, enabling direct MGAR measurements (e.g., Vickers et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023). Vickers et al. (2021, hereafter V21) used LAMOST DR5 and reported a flattening gradient in the thin disk, starting at $0.075\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars and flattening with a rate of $0.003\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$. They also found a slightly positive gradient of $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$ for all ages in the thick disk, consistent with constantly positive gradients of old, cool main-sequence stars ($T_{\mathrm{eff}}<6,000\,\mathrm{K}$) in Yu et al. (2012), which are likely from the thick disk.
@@
Radial migration has been proposed as a major process that alters stellar positions and flattens metallicity gradients (Sellwood & Binney, 2002). This mechanism, sometimes dubbed “churning”, causes significant radial mixing; non-axisymmetric perturbations such as transient spiral arms can induce large-scale stellar migration by exchanging stars across the corotation radius. In contrast, “blurring” refers to orbital heating that increases the amplitude of stellar epicycles without changing their guiding centers. Churning flattens the MGAR as older stars have had more time to migrate, resulting in a flatter radial metallicity distribution.
Recent simulations support the impact of radial migration. Lu et al. (2024) and Ratcliffe et al. (2023) analyzed the correlation between [Fe/H] spreads and stellar birth gradients in mono-age simulated populations, and recovered a time-dependent flattening of birth gradients. Anders et al. (2023) further compared observed MGAR from APOGEE DR17 with modeled ISM gradients evolution from Ratcliffe et al. (2023) to demonstrate the role of migration.
However, the migration interpretation has been challenged by recent FIRE2 simulations (Graf et al., 2024, 2025). They found that present-day MGARs closely follow the stellar birth gradient with minimal influence from churning in these simulations. As gas-phase metal gets mixed before star formation, the role of churning is significantly weakened.
@@
Moreover, Graf et al. (2024) argued that vertical evolution—where turbulent gas settles over time, known as the “upside-down” scenario—may play a more dominant role than inside-out radial growth in shaping Galactic chemo-dynamical history. This suggests the need to consider vertical processes and possible differences in the evolutionary pathways of the thin and thick disks, as also done by V21.
@@
To constrain the MGAR in the MW’s thin and thick disks, we follow an approach similar to V21. We isolate “pure” thin and thick disk populations using both geometric and chemical criteria, and compute MGAR separately for each population. Stellar kinematics and abundances are taken from LAMOST DR8 (Wang et al., 2022a), and stellar ages are calibrated from asteroseismic data provided by the Kepler mission (Wang et al., 2023a), offering an alternative to traditional isochrone fitting.
@@ 2 Data and Method
The Large Area Multi-Object Spectroscopic Telescope (LAMOST), located at Xinglong Station of the National Astronomical Observatory of China, is a 4-meter Schmidt telescope with a 5° field of view. It is one of the most extensive stellar spectroscopic surveys and has released 13 data sets to date. In this work, we use LAMOST Data Release 8 (DR8; Wang et al., 2022a) combined with stellar ages from Wang et al. (2023a), which trained a neural network using asteroseismic data from the LAMOST-Kepler survey.
Our cross-matched sample contains 986,405 unique stars: 688,369 red giant branch (RGB) stars and 298,036 red clump (RC) stars, each with 6D phase-space coordinates, [Fe/H], $\alpha$/Fe], and age measurements. This study treats RGB and RC stars as a single population. We have also tested the MGAR trends separately for RGB and RC stars and found that the qualitative trends remain unchanged.
@@ 2.1 Quality Cuts and Orbital Parameters
To mitigate the impact of orbital blurring, we follow V21 and compute gradients using the Galactocentric radius of the guiding center $R_{\mathrm{g}}$. $R_{\mathrm{g}}$ and the maximum vertical excursion $Z_{\mathrm{max}}$ are calculated using the Galactic potential model MWPotential2014, implemented in the galpy package (Bovy, 2015). We adopt the solar position $R_{0}=8.27\,\mathrm{kpc}$ and circular velocity $v_{0}=236\,\mathrm{km\,s^{-1}}$ (Schönrich & Aumer, 2017). Figure 1 shows the distributions and selection criteria (next Section) used for defining thin/thick disk subsamples.
@@ 2.2 Thin/Thick Disk Subsamples
The MW’s disk can be separated into two components, either geometrically (via vertical excursion) or chemically (via $\alpha$/Fe]–[Fe/H] bimodality). Geometrically, the MW’s stellar density profile can be decomposed into two exponential components (e.g., Gilmore & Reid, 1983). In the $\mathrm{\alpha/Fe}][$$\mathrm{[Fe/H]}$ plane, there are two distinct sequences (e.g., Hayden et al., 2015); such a chemical bimodality reflects differences in nucleosynthetic timescales (Nomoto et al., 2013). The $\alpha$-rich sequence is associated with rapid early star formation dominated by Type II supernovae, which produce more $\alpha$ elements (e.g., $^{\text{16}}$O, $^{\text{20}}$Ne, $^{\text{24}}$Mg, $^{\text{28}}$Si) relative to $^{\text{56}}$Fe. In contrast, Type Ia supernovae, with their longer delay times, enrich more Fe relative to $\alpha$ and contribute more to the $\alpha$-poor (solar-like) sequence (Binney & Tremaine, 2008).
@@ 3 Results
Figure 3 presents the MGAR in the thin and thick disks, compared to the results of V21 (black points). Gradient fits for each mono-age population are shown in detail in Figure 4. Larger uncertainties—particularly for the oldest thin disk stars and the youngest thick disk populations—are primarily due to limited sample sizes (see the age histograms in Figure 2).
We also examine how the MGAR responds to variations in selection criteria and orbital integration configurations in Appendices B and C. Similar to V21, Appendix B shows that the flattening MGAR in the thin disk primarily originates from low-$Z_{\mathrm{max}}$ stars, which may be due to the favored migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014). Appendix C further demonstrates that the choice of Galactic potential $\Phi$ and solar circular velocity $v_{0}$ can affect the numerical values of MGAR in the thin disk, but the overall flattening trend remains robust.
@@ 3.1 Metallicity Gradient in the Thin Disk
The metallicity gradient in the thin disk gradually flattens as stellar age increases and aligns broadly with V21. The gradient of the youngest stars reaches $0.07\,\mathrm{dex\,kpc^{-1}}$ in agreement with V21 and other studies (Chen & Zhao, 2020; Zhang et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023).
A comprehensive comparison of thin disk MGAR measurements using various stellar tracers is given in Wang et al. (2023b). Their Figure 5 highlights a distinctively V-shaped MGAR trend with a dip around $5\sim 7\,\mathrm{Gyr}$ in studies based on main-sequence turn-off (MSTO) stars (e.g., Xiang et al., 2015; Wang et al., 2019), which is absent in our result using RGB and RC stars. V-shapes with a dip around 2 Gyr have been reported using red giant stars from CoRoGEE (Anders et al., 2017) and APOGEE (Anders et al., 2023). The existence and characteristics of this V-shape remain unclear and warrant further investigation with larger samples.
@@ 3.2 Metallicity Gradient in the Thick Disk
The gradients in the thick disk are generally positive across all ages, with a mean value of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$. Interestingly, the gradient increases steadily with decreasing stellar age, reaching a peak of $\sim 0.019\,\mathrm{dex\,kpc^{-1}}$ at $\sim 5.4\,\mathrm{Gyr}$, but flattens at younger ages. This flattening is likely driven by a population of young, metal-poor thick disk stars. As shown in the fourth row of Figure 4, these stars—located at $R_{\mathrm{g}}>10\,\mathrm{kpc}$—contribute to suppressed gradients in the young thick disk panels. Our test confirms that thick disk stars at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ exhibit a negative gradient ($\sim-0.02\,\mathrm{dex\,kpc^{-1}}$). Removing these stars reduces the flattening at young ages and raises the gradient values to $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$, but the overall MGAR trend remains. Overall, our results are consistent with V21, reaffirming the existence of a slightly positive MGAR in the thick disk.
[…]

### v4

@@ Abstract
Galactic disks typically exhibit a negative radial metallicity gradient, indicating faster enrichment in the inner regions. Recent studies report that this gradient becomes flatter with increasing stellar age in the Milky Way’s (MW) thin disk, while the thick disk exhibits a mildly positive gradient across all ages. In this work, we revisit the metallicity gradient–age relation (MGAR) in both the thin and thick disks of the MW, using spectroscopic data from LAMOST DR8 and stellar ages calibrated with asteroseismology. Our results show a steadily flattening MGAR in the thin disk and confirm a positive gradient $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$ in the thick disk. The flattening in the thin disk may be caused by large-scale radial migration induced by transient spiral arms, or by a time-dependent steepening of the interstellar medium (ISM) metallicity gradient as suggested by recent FIRE2 simulations. The positive gradient in the thick disk may reflect early enrichment of the outer regions by strong feedback or starburst-driven outflows in a turbulent, gas-rich proto-disk. These findings suggest distinct chemodynamical evolution paths for the MW’s thin and thick disks and provide valuable constraints for future models of Galactic chemical evolution.
keywords:
Galaxy: abundances – Galaxy: kinematics and dynamics – Galaxy: disc – Galaxy: evolution
pubyear: 2025pagerange: Revisiting the Radial Metallicity Gradient-Age Relation in the Milky Way’s Thin and Thick DisksC
@@ 1 Introduction
A cornerstone of galaxy evolution is the negative radial metallicity gradient, which describes the decrease in stellar metallicity with increasing galactocentric radius. Matteucci and Francois (1989) reproduced this trend using “inside-out” formation models, in which the inner disk forms stars more rapidly and enriches earlier than the outer regions. This scenario naturally results in a negative metallicity gradient for stars formed from the interstellar medium (ISM).
Nowadays, the radial metallicity gradient has become a crucial diagnostic in different classes of models of Galactic evolution. First, chemical evolution models incorporate physical processes such as radial gas flows, star formation histories, and stellar yields to get a realistic metallicity gradient and study the formation of the MW (e.g., Grisoni et al. 2018; Sharda et al. 2021b; see Matteucci 2021 for a review). Second, descriptive models treat the gradient as a parameter to fit observational data and constrain the structure of the Galactic disk (e.g., Schönrich and Binney, 2009; Sanders and Binney, 2015; Frankel et al., 2020; Sharma et al., 2021). Finally, hybrid models combine chemical evolution with numerical simulations, as in Minchev et al. (2013) and Johnson et al. (2021). Utilizing the hybrid framework of Minchev et al. (2013), Minchev et al. (2018) inferred stellar birth radius and concluded that the ISM gradient flattens over time as the outer disk becomes progressively enriched and approaches the central metallicity.
@@
The pioneering study on the MGAR in our MW by Mayor (1976) measured a mean radial metallicity gradient of $0.05\pm 0.01\,\mathrm{dex\,kpc^{-1}}$, which steepens to $0.10\pm 0.02\,\mathrm{dex\,kpc^{-1}}$ for younger stars. Using effective temperature ($T_{\mathrm{eff}}$) of main-sequence stars as a proxy for age, Yu et al. (2012) found that the gradient steepens with increasing $T_{\mathrm{eff}}$ (typically associated with decreasing age), but shows no trend for stars with $T_{\mathrm{eff}}<6,000\,\mathrm{K}$ (i.e., those with main-sequence lifetimes $\gtrsim 10$ Gyr). These early findings hinted at a flattening gradient for increasingly older stars in the MW’s thin disk.
@@
Large-scale spectroscopic surveys have greatly facilitated precise age estimates, enabling direct MGAR measurements (e.g., Vickers et al., 2021; Wang et al., 2023b; Willett et al., 2023; Ratcliffe et al., 2023). Vickers et al. (2021, hereafter V21) used LAMOST DR5 and reported a flattening gradient in the thin disk, starting at $0.075\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars and flattening with a rate of $0.003\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$. Willett et al. (2023) used asteroseismic age and adopted a hierarchical Bayesian model for the gradient fitting. They found a similar flattening trend as V21 in the thin disk. In the thick disk, V21 found a slightly positive gradient of $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$ across all age bins, consistent with constantly positive gradients of old, cool main-sequence stars ($T_{\mathrm{eff}}<6,000\,\mathrm{K}$) in Yu et al. (2012), which are likely from the thick disk.
@@
Radial migration has been proposed as a major process that alters stellar positions and flattens metallicity gradients (Sellwood and Binney, 2002). This mechanism, sometimes dubbed “churning”, causes significant radial mixing; non-axisymmetric perturbations such as transient spiral arms can induce large-scale stellar migration by exchanging stars across the corotation radius. In contrast, “blurring” refers to orbital heating that increases the amplitude of stellar epicycles without changing their guiding centers. Churning flattens the MGAR as older stars have had more time to migrate, resulting in a flatter radial metallicity distribution.
The metallicity gradients at stellar birth provided by the recent method [^1] https://github.com/BridgetRatcliffe/Rbirth for estimating the stellar birth radius ($R_{\mathrm{birth}}$) further supports the significance of radial migration. Building on the empirical framework of Minchev et al. (2018), a more systematic approach to infer $R_{\mathrm{birth}}$ from stellar ages and metallicities has been developed in the follow-up studies (Lu et al., 2024; Ratcliffe et al., 2023, 2025b, 2025a). In these works, the evolution of the stellar birth metallicity gradient is approximated as
$\nabla\mathrm{[Fe/H]}(\tau)\approx\frac{\Delta\mathrm{[Fe/H]}(\tau)}{\Delta R(\tau)}\propto\frac{\mathrm{Range[Fe/H](age)}}{\Delta R(\tau)},$ (1)
where $\tau$ is the lookback time and equals stellar age. A linear anti-correlation between $\nabla\mathrm{[Fe/H]}(\tau)$ and $\mathrm{Range[Fe/H](age)}$—validated with NIHAO-UHD (Buck et al., 2020) and HESTIA (Libeskind et al., 2020) simulations by Lu et al. (2024)—then provides an estimate of the birth gradient $\nabla\mathrm{[Fe/H]}(\tau)$ (with later refinements correcting $\Delta R(\tau)$ and accounting for selection effects; see Ratcliffe et al. 2025b, a). Finally, the birth radius is obtained as
$R_{\mathrm{birth}}(\mathrm{age},\mathrm{[Fe/H]})=\frac{\mathrm{[Fe/H]}-\mathrm{[Fe/H]}(R=0,\tau)}{\nabla\mathrm{[Fe/H]}(\tau)},$ (2)
given an estimate of $\mathrm{[Fe/H]}(R=0,\tau)$.
As a by-product of this method, the reconstructed stellar birth gradients $\nabla\mathrm{[Fe/H]}(\tau)$ become progressively flatter with decreasing lookback time/age. Anders et al. (2023) further compared APOGEE MGAR measurements with the flattening birth-gradient evolution of Ratcliffe et al. (2023), demonstrating how radial migration could explain the difference between birth gradient evolution and observed MGAR trend.
On the simulation side, models with smooth/bursty stellar feedback predict qualitatively different birth-gradient evolutions with redshift. As summarized in Garcia et al. (2025b), simulations with smooth feedback (such as EAGLE, Illustris, IllustrisTNG, and SIMBA) employ subgrid treatments of the ISM that produce relatively smooth metal redistribution and predict strong negative gradients flattening with time. In contrast, bursty-feedback simulations (e.g., FIRE2, SPICE Bursty, and Thesan Zoom) yield relatively shallower birth gradients at all epochs and better match current observations at high redshifts (Garcia et al., 2025a).
In particular, recent FIRE2 simulations strongly challenge the migration-based interpretation of the MGAR (Graf et al., 2024, 2025). In these models, the present-day MGAR closely tracks the stellar birth gradients, implying minimal stellar radial migration in the simulations. And as an alternative to stellar migration, gas mixing (diffusion, advection, and turbulence) efficiently mixes gas-phase metals before forming new stars.
@@
Moreover, Graf et al. (2024) argued that vertical evolution, where turbulent gas gradually settles over time (“upside-down” disk formation), may play a more important role than inside-out radial growth in shaping the Galaxy’s chemo-dynamical history. This highlights the need to account for vertical processes and potentially distinct evolutionary pathways of the thin and thick disks, as also emphasized by V21.
@@
To constrain the MGAR in the MW’s thin and thick disks, we follow an approach similar to V21. We isolate “pure” thin and thick disk populations using both geometric and chemical criteria, and compute MGAR separately for each population. Stellar kinematics and abundances are taken from LAMOST DR8 (Wang et al., 2022a), and stellar ages are calibrated from asteroseismic data provided by the Kepler mission (Gilliland et al., 2010; Borucki et al., 2010; Wang et al., 2023a), offering an alternative to traditional isochrone fitting.
@@ 2 Data and Method
The Large Area Multi-Object Spectroscopic Telescope (LAMOST, Deng et al., 2012; Zhao et al., 2012; Liu et al., 2014), located at Xinglong Station of the National Astronomical Observatory of China, is a 4-meter Schmidt telescope with a 5° field of view. It is one of the most extensive stellar spectroscopic surveys and has released 13 data sets to date. In this work, we use LAMOST Data Release 8 (DR8; Wang et al., 2022a) combined with stellar ages from Wang et al. (2023a), which trained a neural network using asteroseismic data from the LAMOST-Kepler common stars.
Our cross-matched sample comprises 986,405 unique stars, including 688,369 red giant branch (RGB) stars and 298,036 red clump (RC) stars, each with 6D phase-space coordinates, [Fe/H], $\alpha$/Fe], and age measurements. This study treats RGB and RC stars as a single population. We have also tested the MGAR trends separately for RGB and RC stars and found that the qualitative trends remain unchanged.
@@ 2.1 Quality cuts and orbital parameters
To mitigate the impact of orbital blurring, we follow V21 and compute gradients using the Galactocentric radius of the guiding center $R_{\mathrm{g}}$. $R_{\mathrm{g}}$ and the maximum vertical excursion $Z_{\mathrm{max}}$ are calculated using the Galactic potential model MWPotential2014, implemented in the galpy package (Bovy, 2015). We adopt the solar position $R_{0}=8.27\,\mathrm{kpc}$ and circular velocity $v_{0}=236\,\mathrm{km\,s^{-1}}$ (Schönrich and Aumer, 2017). Figure 1 shows the distributions and selection criteria (next Section) used for defining thin/thick disk subsamples.
@@ 2.2 Thin/thick disk subsamples
The MW’s disk can be separated into two components, either geometrically (via vertical excursion) or chemically (via $\alpha$/Fe]–[Fe/H] bimodality). Geometrically, the MW’s stellar density profile can be decomposed into two exponential components (e.g., Gilmore and Reid, 1983). In the $\mathrm{\alpha/Fe}][$$\mathrm{[Fe/H]}$ plane, there are two distinct sequences (e.g., Hayden et al., 2015; Imig et al., 2023); such a chemical bimodality reflects differences in nucleosynthetic timescales (Nomoto et al., 2013). The $\alpha$-rich sequence is associated with rapid early star formation dominated by Type II supernovae, which produce more $\alpha$ elements (e.g., $^{\text{16}}$O, $^{\text{20}}$Ne, $^{\text{24}}$Mg, $^{\text{28}}$Si) relative to $^{\text{56}}$Fe. In contrast, Type Ia supernovae, with their longer delay times, enrich more Fe relative to $\alpha$ and contribute more to the $\alpha$-poor (solar-like) sequence (Binney and Tremaine, 2008).
@@ 3 Results
Figure 3 presents the MGAR in the thin and thick disks, compared to the results of V21 (black points). Gradient fits for each mono-age population are shown in detail in Figure 4. Larger uncertainties for the oldest thin disk stars and the youngest thick disk populations are primarily due to limited sample sizes (see the age histograms in Figure 2).
We also examine how the MGAR responds to variations in selection criteria and orbital integration configurations in Appendices B and C. Similar to V21, Appendix B shows that the flattening MGAR in the thin disk primarily originates from low-$Z_{\mathrm{max}}$ stars, which may be due to the favored migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014). For the intermediate $Z_{\mathrm{max}}$, we observe a slightly non-monotonic MGAR, similar to V21 and Willett et al. (2023). Appendix C further demonstrates that the choice of Galactic potential $\Phi$ and solar circular velocity $v_{0}$ can affect the numerical values of MGAR in the thin disk, but the overall flattening trend remains robust.
@@ 3.1 Metallicity gradient in the thin disk
The metallicity gradient in the thin disk gradually flattens as stellar age increases and aligns broadly with V21 and Willett et al. (2023). The gradient of the youngest stars reaches $0.07\,\mathrm{dex\,kpc^{-1}}$ in agreement with V21 and other studies (Chen and Zhao, 2020; Zhang et al., 2021; Wang et al., 2023b; Willett et al., 2023; Ratcliffe et al., 2023; Sun et al., 2024).
A comprehensive comparison of thin disk MGAR measurements using various stellar tracers is given in Wang et al. (2023b). Their Figure 5 highlights a distinctively V-shaped MGAR trend with a dip around $5\sim 7\,\mathrm{Gyr}$ in studies based on main-sequence turn-off (MSTO) stars (e.g., Xiang et al., 2015; Wang et al., 2019), which is absent in our result using RGB and RC stars. V-shapes with a dip around 2 Gyr have been reported using red giant stars from CoRoGEE (Anders et al., 2017) and APOGEE (Anders et al., 2023). It is worth noting that the V-shapes seem to appear more frequently when the gradient is calculated with respect to $R_{\mathrm{present}}$ (Xiang et al., 2015; Anders et al., 2017; Wang et al., 2019) rather than $R_{\mathrm{g}}$ (V21; Willett et al., 2023), though the V-shape of MGAR in Anders et al. (2023) is not sensitive to either radius. The existence and characteristics of these V-shapes remain unclear and warrant further investigation with larger samples.
[…]
@@ 3.2 Metallicity gradient in the thick disk
The gradients in the thick disk are generally positive across all ages, with a mean value of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$. Li and Zhao (2017) reported a higher value of $0.035\,\mathrm{dex\,kpc^{-1}}$ using 2,035 thick-disk giant stars from LAMOST DR3. Recently, Hu et al. (2023) measured the metallicity gradient with respect to the angular momentum $L_{z}$ in their chemically thick (high-$\alpha$ metal-poor) disk: $1.2\times 10^{-3}\,\mathrm{dex\,kpc^{-1}\,km^{-1}\,s}$, which corresponds to $\mathrm{d[M/H]}/\mathrm{d}R_{\mathrm{g}}=0.28\,\mathrm{dex\,kpc^{-1}}$ and is significantly larger than our findings here. Sun et al. (2024) found a positive gradient weaker than $\sim 0.01\,\mathrm{dex\,kpc^{-1}}$ in the high-$\alpha$ disk across different $|Z|$ bins.
Interestingly, our gradient increases steadily with decreasing stellar age, reaching a peak of $\sim 0.019\,\mathrm{dex\,kpc^{-1}}$ at $\sim 5.4\,\mathrm{Gyr}$, but flattens at younger ages. This flattening is likely driven by a population of young, metal-poor thick disk stars. As shown in the fourth row of Figure 4, the stars located at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ help to suppress gradients in the young thick disk panels. Our test confirms that thick disk stars at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ exhibit a negative gradient ($\sim-0.02\,\mathrm{dex\,kpc^{-1}}$). Removing these stars reduces the flattening at young ages and raises the gradient values to $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$, but the overall MGAR trend remains. Overall, our results are consistent with V21, reaffirming the existence of a slightly positive MGAR in the thick disk.